Under conditions of constant stress, Newtonian flow is observed for the Maxwell model Starting with the Maxwell’s equation for shear rate, prove that this is true.
Answer #1
  • A Newtonian fluid is independent of both time and rate of shear.
  • Newtonian model describes an idealized type of flow. In many systems, a material may exhibit Newtonian behaviour over a wide range of shear rate, but may, surprisingly, demonstrate non-Newtonian behaviour outside of that range.
  • For Such a conditions equation using Maxwell Model are describe below.
  • The Maxwell model is a two-element model consisting of a linear spring element and a linear viscous dashpot element connected in series, as shown in fig.

  • The stress-strain relations of spring and dashpot respectively are:

    σ=Eɛ2

    σ=η x dɛ/dt

    Because both elements are connected in series, the total strain is:

    ɛ=ɛ1+ɛ2  

    By eliminating ɛ1 and ɛ2, the following constitutive equation for the Maxwell model is obtained:

    dɛ/dt=(1/E)(dσ/dt) +(σ/η)

    The response of the model to various stress or strain conditions can be obtained after applying integration together with appropriate initial conditions.

  • applying a constant stress σ = σo at t = 0, the following response is obtained:

    ɛ(t) =(σ/E) +(σ/η)t

  • From above equation it is prove that at constant stress for Newtonian flow is observed for the Maxwell Model starting with the Maxwell' equation for shear rate.